OpenStudy (anonymous):

Which expression is equivalent to (5^3)^−2

OpenStudy (anonymous):

http://akk.li/pics/anne.jpg

OpenStudy (anonymous):

ur too mean

OpenStudy (anonymous):

heehee

OpenStudy (cwrw238):

use the rule (a^m)^n = a^(mn) and a^(-n) = 1 / a^n

OpenStudy (cwrw238):

aaah!!

OpenStudy (cwrw238):

lol

OpenStudy (anonymous):

i just saw that answer because wingspanhead dorothy up there made me close the computer and hide under thhe covers